One-year Treasury securities yield 5.5 percent. The market anticipates that 1 year from now, 1-year Treasury securities will yield 6.5 percent. If the pure expectations theory is correct, what is the yield today for 2-year Treasury securities? 6.5% 6.0% 75% 3.5% 4.5%
Answer #1

Here 1 year treasury yield r1= 5.5%

1 year foward rate f2 = 6.5%

2 years treasury rate r2 =?

As per epectation theroy below is the formula for 2 year spot rate:

(1+r2)^2 = (1+r1)*(1+f2)

(1+r2)^2 = (1+5.5%)(1+6.5%)

= 1.055*1.065

(1+r2)^2 = 1.123575

(1+r2) = 1.059988

r2 = 6.00%

So second option is correct here.

Thanks